A rectangular garden has a length of 5ft and a width of 8 ft. The length is increased by 2 feet. By how many square feet does the area increase?
Answer by drj(1380)

A rectangular garden has a length of 5ft and a width of 8 ft. The length is increased by 2 feet. By how many square feet does the area increase?

Step 1. Area of a rectangle is given as A=length * width

Step 2. Then, A1=5*8=40 square feet of original area of rectangular garden.

Step 3. Then, A2=(5+2)*8=56 square feet since the length is increased by 2. feet.

Step 4. Difference = A2 - A1 = 56 - 40 = 16 square feet.

Step 5. ANSWER: The area increased by 16 square feet.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J